40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11813     Entry time: Wed Nov 25 22:37:12 2015
Author: yutaro 
Type: Update 
Category: LSC 
Subject: removal of Gautam's cable in 1Y2 and restoration of POYDC 

[yutaro, Koji]

We disconnected the cable that was connected to CH5 of the whitening filter in 1Y2, then connected POYDC cable to there (CH5). This channel is where POYDC used to connect.

Then we turned on the whitening filter for POYDC (C1:LSC-POYDC FM1) and changed the gain of analog whitening filter for POYDC from 0 dB to 39 dB (C1:LSC-POYDC_WhiteGain).

ELOG V3.1.3-